# Elk: Exploring the Efficiency of Inter-core Connected AI Chips with Deep Learning Compiler Techniques

### v1

@@ Abstract.
copyright: nonejournalyear: 2025conference: The 58th IEEE/ACM International Symposium on Microarchitecture; October 18–22, 2025; Seoul, Korea
@@ 1. Introduction
To meet the ever-increasing compute demand of deep learning (DL) like large language models (LLMs) (Narayanan et al., 2021; Kwon et al., 2023), various AI chips have been developed (NVIDIA H100 Tensor Core GPU, [n. d.]; Knowles, 2021; Prabhakar et al., 2024; Meta, 2024; Tenstorrent, 2023). A typical AI chip employs many parallel cores to scale computing throughput. Each core has its local SRAM as a scratchpad memory. To exploit this parallelism, the DL compiler partitions a tensor operator (e.g., BatchMatMul in attention and MatMul in FFN (Vaswani et al., 2023)) into tiles and maps each tile to a core. Since the on-chip SRAM size is limited, AI chips can employ off-chip memories (e.g., HBM) to provide larger capacity and accommodate the model parameters of larger DL models.
However, the off-chip memory bandwidth scales much slower than compute performance, and cannot meet the growing demand of large models. To alleviate the bandwidth bottleneck, inter-core connected AI (ICCA) chips were proposed. They enable inter-core links that allow one core to directly access data from other cores’ SRAM, as shown in Figure 1. A typical ICCA chip example is Graphcore IPU (Knowles, 2021). It has 1472 cores, each core has 624KB local SRAM and can access another core’s SRAM at 5.5GB/s. This aggregates to an 896MB on-chip memory with 8TB/s all-to-all data exchange bandwidth. The large on-chip space and high memory bandwidth present a promising way to break the memory wall for DL workloads (e.g., compared to an A100 GPU with 60MB total cache size and 2TB/s HBM bandwidth). With these advantages, inter-core interconnect has been employed by many AI chips today, such as Graphcore IPU (Knowles, 2021), SambaNova SN40 (Prabhakar et al., 2024), Cerebras WSE (Lie, 2021), Meta MTIA (Meta, 2024), and NVIDIA’s H100 GPU (NVIDIA H100 Tensor Core GPU, [n. d.]).
@@
The inter-core interconnect connects all cores’ local SRAM into a distributed memory space that can be managed by software (i.e., compiler), leading to new parallel execution models for DL workloads (§$\S$2.2). In conventional accelerators without inter-core connections, all cores execute independently with their local SRAM, and a separate global SRAM shared by all cores simultaneously handles all off-chip data loading. On ICCA chips, the software can manually manage data sharing among cores without needing a global SRAM. Also, the distributed nature of ICCA chip’s on-chip SRAM allows its size to further scale, so it can store multiple tensor operators. Thus, when executing a current operator, the chip can simultaneously preload future operators’ data from off-chip memory to SRAM. However, this requires each core’s local SRAM to enable double buffering between execution and preload, resulting in significant memory footprint overhead.
@@
First, to overlap computing and off-chip loading, the DL compiler needs to decide how much on-chip memory space to allocate for per-core execution and for buffering preloaded data from off-chip memory. A larger space for execution (i.e., execution space) allows larger per-core tile size, reduces inter-core communication traffic, and improves compute efficiency. A larger space for preload (i.e., preload space) improves off-chip memory bandwidth utilization. This leads to on-chip memory capacity contention ( 1 in Figure 2). Second, the on-chip interconnect links all cores and HBM controllers, and its bandwidth is shared between inter-core data exchange and HBM-to-core data loading. This leads to interconnect bandwidth contention ( 2). Third, the per-core SRAM must feed data to the local computation pipeline and serve data to other cores via the interconnect. The concurrent SRAM accesses will lead to memory access contention ( 3).
@@
Figure 2. Resource contentions on ICCA chip with HBM.
In this paper, we present Elk, a DL compiler framework to maximize the efficiency of ICCA chips by jointly optimizing all the three performance factors. Elk formalizes these factors into a global trade-off space, based on the insight that these factors can be transferred into configurable compiler parameters (§$\S$3), and the correlation between these parameters can reflect their performance trade-offs.
@@
  • With Elk and the ICCA chip simulator, we enable design space exploration of ICCA chips and present our insights in §$\S$6.4. We will open source our codebase to the community.
@@ 2.1. Architecture of the ICCA Chip
[…]
Scale the ICCA chip with HBM. To serve large models whose sizes exceed the on-chip capacity, we can scale the memory capacity with off-chip memory modules like high bandwidth memory (HBM) (High Bandwidth Memory, [n. d.]). Many ICCA chips already integrate off-chip memory (Prabhakar et al., 2024; Meta, 2024). As shown in Figure 1, they attach HBM controllers to the on-chip interconnect, so each controller can directly send data to each core similar to how the cores send data to each other. To access HBM, cores communicate with HBM controllers via the interconnect. The HBM controller coalesces the memory requests from cores, loads data from HBM, and sends data to cores.
@@ 2.3. Challenges of Using ICCA Chip with HBM
To maximize the ICCA chip performance, we must ( 1) allow faster per-core execution, ( 2) utilize more HBM bandwidth to preload required data on time, and ( 3) reduce the inter-core data sharing overhead. However, it is difficult to maximize all three performance metrics simultaneously, as they have conflicting resource demands.
On-chip memory space contention. We cannot 1 maximize per-core execution performance and 2 HBM bandwidth utilization at the same time, due to on-chip memory space contention. As shown by 1 in Figure 2, each core reserves an execution space for the currently executing operator and a preload space for the preloaded operators. To speed up per-core execution, a larger execution space is required (see §3.1 and Figure 5). To prevent HBM underutilization, a larger preload space is required (see §3.2 and Figure 6). With limited on-chip memory, we cannot expand both spaces.
Interconnect bandwidth contention. We cannot 2 maximize HBM bandwidth utilization and 3 minimize inter-core data sharing overhead at the same time, due to the interconnect bandwidth contention. As shown by 2 in Figure 2, the on-chip interconnect carries both core-to-core traffic for inter-core data sharing and HBM controller-to-core traffic for preloading. When both traffic flows are heavy, the interconnect will be congested (see §3.3 and Figure 8).
Memory access contention. We cannot 1 maximize per-core execution performance and 3 minimize inter-core data sharing overhead at the same time, due to the memory access contention. As shown by 3 in Figure 2, each core’s local memory is simultaneously accessed by the core itself for computing a tile, and by other cores for inter-core data sharing. For example on IPU, each core reads its local memory at full speed (128 bits/cycle (Graphcore, 2022)) when executing DL operators like MatMul, any other accesses will pause the execution. Upon contention, tile execution on this core reads data from local memory at slower speed, or even pauses entirely. The remote cores may also suffer from degraded SRAM bandwidth.
@@ 3. Performance Tradeoffs in Elk
As discussed in §$\S$2.3, to efficiently use ICCA chips with HBM, we must trade-off multiple performance factors. We summarize how each performance factor is mapped to a compiler decision in Figure 4. First, increasing per-core execution space enables faster per-core execution with a larger tile size (§3.1). Second, increasing number of preloaded operators can better overlap on-chip execution and off-chip HBM load, improving HBM bandwidth utilization (§3.2). Third, increasing preload space for each preloaded operator reduces the inter-core data exchange overhead and the memory access contention, since the shared data can be duplicated on cores in advance to reduce the overhead of on-demand accesses to other cores (§3.3). We validate the insights with experiments on our ICCA chip emulator (see implementation in §$\S$5) as follows.
@@ 3.1. Larger Execution Space Enables Faster Per-core Execution
While existing compilers focus on achieving higher on-chip execution performance using a given execution space size, however, they cannot find a proper execution space size by arbitrating the memory space contention in §2.3. Moreover, operators from the same model have diverse memory vs. time correlations. Thus, we also need to adjust the execution space size based on each operator’s performance characteristic, rather than allocating a fixed-sized execution space throughout the model execution.
@@ 3.3. Larger Per-Operator Preload Space Reduces Inter-core Data Access Volume
Figure 7 shows that expanding preload space reduces the inter-core bandwidth demand. For each operator, we pick the fastest execution plan that fits in a given execution space size [^1] For each run, we use the optimal execution space size that gives the smallest total inference latency. See the description of the Static setup in §6.1.. MinPreload lets each core access all shared data from other cores at execution time, which requires the minimum preload space. MaxPreload lets HBM controllers broadcast as much shared data as possible at preload time, which requires the largest preload space. We profile the inter-core bandwidth demand (inter-core transfer volumeper-core execution $\frac{\text{inter-core transfer volume}}{\text{per-core execution time}}$) of each core. MaxPreload significantly reduces the inter-core traffic.
Although more broadcasts on preload increase the HBM controller-to-core traffic, the preload traffic can be opportunistically interleaved with ongoing inter-core traffic to reduce contention. Figure 8 shows how each core’s total interconnect bandwidth demand (defined as inter-core transfer volumeper-core execution time+HBM-to-core transfer volumeHBM load $\vspace{-0.2em}\frac{\text{inter-core transfer volume}}{\text{per-core execution time}}+\frac{\text{HBM-to-core transfer volume}}{\text{HBM load time}}\vspace{0.1em}$) varies over time. Purely relying on inter-core transfer fluctuates the traffic pressure drastically, causing interconnect underutilization or congestion. More broadcasts at preload time reduce fluctuation by spreading the traffic across preload and execution times.
Figure 8. The total per-core interconnect bandwidth demand.
@@ 4.2. Two-Level Inductive Operator Scheduling
The base case of induction is trivial, as the last operator has no succeeding operators to preload (i.e., preload number is always 0). For the inductive step, we show an example in Figure 10. In Figure 10 (a), Elk has finished scheduling all operators after Op5. Then, Elk schedules the execution of Op5 in Figure 10 (b). Elk enumerates all possible preload numbers for Op5. For each preload number, Elk invokes the cost-aware on-chip memory allocation algorithm (§$\S$4.3) to determine the execution/preload space sizes for the involved operators and the estimated execution time of Op5.
For example, preload number 0 means we do not overlap Op5’s execution with any preload. The execution time of Op5 is minimized, but the overall execution time from Op5 to the end of the model is sub-optimal. For preload numbers 1 and 2, Elk overlaps the execution of Op5 with the preload of Op6, or the preloads of both Op6 and Op7. Though the execution time of Op5 is longer, the overall execution times are better than preload number 0. As preload number 1 yields the lowest current-to-end time, Elk selects it for Op5.
@@
Our algorithm has $O(KN)$ complexity as we iterate through $N$ operators and at most $K$ preload numbers per operator. The algorithm provably finds the end-to-end plan with the shortest total time, assuming it can obtain the optimal execution time for each preload number. Lemma 4.1 and Theorem 4.2 formalizes the algorithm (see Appendix A for proofs).
@@
[…]
@@ 4.3. Cost-Aware On-chip Memory Allocation
First, there are two types of intra-operator tradeoffs: (1) For the currently executing operator, Elk trades memory space for execution time (§$\S$3.1). (2) For each preloaded operator, Elk trades memory space for this operator’s inter-core data exchange overhead (§$\S$3.3). Second, there is an inter-operator tradeoff: as operators have different memory-time tradeoffs, we allocate more memory to operators that benefit more from a larger execution/preload space.
@@
Cost model for execution time. As DL workloads have predictable execution patterns (Feng et al., 2023; Miao et al., 2022; Zhu et al., 2022; Liu et al., 2024; Agrawal et al., 2024), Elk uses an accurate cost model to quickly estimate the performance of per-core execution and inter-core transfer. For each operator type (e.g., MatMul), we randomly generate tiles with varied shapes, and run each tile using one core on the target device. Then, we fit a linear tree model (Cerliani, 2022) using the tile shapes as inputs and the profiled execution times as outputs. For inter-core transfer, we fit a model for each network link using transfer volumes as inputs and transfer times as outputs. For each partition plan, Elk determines the tile-to-core mapping and orchestrates the inter-core transfer (e.g., the source/destination cores and intermediate hops of each transfer, see §$\S$5). Elk uses the per-link cost model and the communication pattern to estimate the total transfer time. Figure 12 shows that Elk can accurately predict the execution and transfer times of an IPU chip. Elk can use different cost models (Miao et al., 2022; Zhu et al., 2022; Liu et al., 2024; Agrawal et al., 2024) for different hardware platforms.
Intra-operator tradeoff for preloading (Tradeoff 2 and 3 in Figure 11). For each preloaded operator, its partition plan is already decided in a previous step of the inductive operator scheduling (§$\S$4.2). This execute-state plan is chosen for execution speed, which may use more memory space. As the operator is not currently executing, Elk assigns a memory-efficient preload-state plan. To start execution, a data distribution phase transforms the operator from preload- to execute-state by distributing the required data via the interconnect (e.g., Figure 3 (c)). It saves this operator’s preload space at the cost of extra inter-core data exchange overhead, compared to broadcasting the required data at preload time following the execute-state plan (e.g., Figure 3 (b)).
Each execute-state plan may have many preload-state plans, by configuring how much data is broadcasted on preload. On preload, if 4 cores share a data piece, we can evenly split it into 1, 2, or 4 chunks, and broadcast each chunk to 4, 2, or 1 cores. Each core receives $1$, $\frac{1}{2}$, or $\frac{1}{4}$ of the data on preload (this decides preload space size), and fetches the rest 00, $\frac{1}{2}$, or $\frac{3}{4}$ on data distribution. Elk finds the Pareto-optimal preload-state plans of each preloaded operator, by estimating their preload space sizes and data distribution times.
@@
For each search step, Elk examines the next plan with a smaller memory footprint along the Pareto-optimal curve for each operator. Elk selects the most “cost-effective” operator whose next plan has the largest ratio Δ=reduced space sizeincreased $\Delta=\frac{\text{reduced space size}}{\text{increased time}}$ compared to the currently selected plan. For example, in Figure 11, Op5 is the executing operator. Op6 and Op7 are preloaded operators. Elk updates the current plan for Op7 and proceed to the next search step. Elk stops when the total memory requirement does not exceed the available capacity.
In the worst case, Elk needs to examine all Pareto-optimal plans for all operators. Hence, it has $O(PK)$-time complexity for $K$ operators to fit on-chip and $P$ plans per operator. Combined with §$\S$4.2, the complexity is $O(PK^{2}N)$ for $N$ operators ($K$ is also the number of possible preload numbers of each operator).

### v2

@@ Abstract.
Deep Learning Compiler, Inter-Core Connected AI Chip, ML Accelerator, Distributed On-chip Memory
journalyear: 2025copyright: ccconference: 58th IEEE/ACM International Symposium on Microarchitecture; October 18-22, 2025; Seoul, Republic of Koreabooktitle: 58th IEEE/ACM International Symposium on Microarchitecture (MICRO’25), October 18-22, 2025, Seoul, Republic of Koreadoi: 10.1145/3725843.3756064isbn: 979-8-4007-1573-0/2025/10ccs: Software and its engineering Compilersccs: Hardware Emerging architecturesccs: Computer systems organization Parallel architectures
@@ 1. Introduction
To meet the ever-increasing compute demand of deep learning (DL) like large language models (LLMs) (Narayanan et al., 2021; Kwon et al., 2023), various AI chips have been developed (NVIDIA H100 Tensor Core GPU, 2024; Knowles, 2021; Prabhakar et al., 2024; Meta, 2024; Tenstorrent, 2023). A typical AI chip employs many parallel cores to scale computing throughput. Each core has its local SRAM as a scratchpad memory. To exploit this parallelism, the DL compiler partitions a tensor operator (e.g., BatchMatMul in attention and MatMul in FFN (Vaswani et al., 2023)) into tiles and maps each tile to a core. Since the on-chip SRAM size is limited, AI chips can employ off-chip memories (e.g., HBM) to provide larger capacity and accommodate the model parameters of larger DL models.
However, the off-chip memory bandwidth scales much slower than compute performance, and cannot meet the growing demand of large models. To alleviate the bandwidth bottleneck, inter-core connected AI (ICCA) chips were proposed. They enable inter-core links that allow one core to directly access data from other cores’ SRAM, as shown in Figure 1. A typical ICCA chip example is Graphcore IPU (Knowles, 2021). It has 1472 cores, each core has 624KB local SRAM and can access another core’s SRAM at 5.5GB/s. This aggregates to an 896MB on-chip memory with 8TB/s all-to-all data exchange bandwidth. The large on-chip space and high memory bandwidth present a promising way to break the memory wall for DL workloads (e.g., compared to an A100 GPU with 60MB total cache size and 2TB/s HBM bandwidth). With these advantages, inter-core interconnect has been employed by many AI chips today, such as Graphcore IPU (Knowles, 2021), SambaNova SN40 (Prabhakar et al., 2024), Cerebras WSE (Lie, 2021), Meta MTIA (Meta, 2024), and NVIDIA’s H100 GPU (NVIDIA H100 Tensor Core GPU, 2024).
@@
The inter-core interconnect connects all cores’ local SRAM into a distributed memory space that can be managed by software (i.e., compiler), leading to new parallel execution models for DL workloads ($\S$2.2). In conventional accelerators without inter-core connections, all cores execute independently with their local SRAM, and a separate global SRAM shared by all cores simultaneously handles all off-chip data loading. On ICCA chips, the software can manually manage data sharing among cores without needing a global SRAM. Also, the distributed nature of ICCA chip’s on-chip SRAM allows its size to further scale, so it can store multiple tensor operators. Thus, when executing a current operator, the chip can simultaneously preload future operators’ data from off-chip memory to SRAM. However, this requires each core’s local SRAM to enable double buffering between execution and preload, resulting in significant memory footprint overhead.
@@
First, to overlap computing and off-chip loading, the DL compiler needs to decide how much on-chip memory space to allocate for per-core execution and for buffering preloaded data from off-chip memory. A larger space for execution (i.e., execution space) allows larger per-core tile size, reduces inter-core communication traffic, and improves compute efficiency. A larger space for preload (i.e., preload space) improves off-chip memory bandwidth utilization. This leads to on-chip memory capacity contention ( in Figure 2). Second, the on-chip interconnect links all cores and HBM controllers, and its bandwidth is shared between inter-core data exchange and HBM-to-core data loading. This leads to interconnect bandwidth contention ( ). Third, the per-core SRAM must feed data to the local computation pipeline and serve data to other cores via the interconnect. The concurrent SRAM accesses will lead to memory access contention ( ).
Figure 2. Resource contentions on ICCA chip with HBM.
@@
In this paper, we present Elk, a DL compiler framework to maximize the efficiency of ICCA chips by jointly optimizing all the three performance factors. Elk formalizes these factors into a global trade-off space, based on the insight that these factors can be transferred into configurable compiler parameters ($\S$3), and the correlation between these parameters can reflect their performance trade-offs.
@@
  • With Elk and the ICCA chip simulator, we enable design space exploration of ICCA chips and present our insights in $\S$6.4. We will open source our codebase to the community.
@@ 2.1. Architecture of the ICCA Chip
To facilitate the introduction of the ICCA chip as shown in Figure 1, we use Graphcore IPU MK2 (Knowles, 2021) as an example. An IPU chip has 1472 cores that execute independently in parallel. Each core has 624KB local scratchpad memory, adding up to 896MB of total on-chip memory. All cores are interconnected with high-bandwidth low-latency links. Each core can access any other core’s local memory at 5.5GB/s, delivering an aggregated inter-core all-to-all bandwidth of $1472\times 5.5\text{GB/s}\approx 8\text{TB/s}$ (Jia et al., 2020). The large on-chip memory improves on-chip data reuse by storing more operators or even an entire model. The all-to-all interconnect allows each core to independently access on-chip data at high bandwidth. If multiple cores receive/send different data from/to the same core, the interconnect sequentially serves each data transfer at full bandwidth. In addition to IPU, other ICCA chips, such as SambaNova SN40L (Prabhakar et al., 2024) and Tenstorrent (Tenstorrent, 2023), feature a mesh-based on-chip interconnect. In general, the ICCA chip architecture enables scalable performance and alleviates the memory bandwidth bottleneck for serving memory-intensive DL workloads like LLMs.
Scale the ICCA chip with HBM. To serve large models whose sizes exceed the on-chip capacity, we can scale the memory capacity with off-chip memory modules like high bandwidth memory (HBM) (Micron, 2024). Many ICCA chips already integrate off-chip memory (Prabhakar et al., 2024; Meta, 2024). As shown in Figure 1, they attach HBM controllers to the on-chip interconnect, so each controller can directly send data to each core similar to how the cores send data to each other. To access HBM, cores communicate with HBM controllers via the interconnect. The HBM controller coalesces the memory requests from cores, loads data from HBM, and sends data to cores.
@@ 2.3. Challenges of Using ICCA Chip with HBM
To maximize the ICCA chip performance, we must (1) allow faster per-core execution, (2) utilize more HBM bandwidth to preload required data on time, and (3) reduce the inter-core data sharing overhead. However, it is difficult to maximize all three performance metrics simultaneously, as they have conflicting resource demands.
On-chip memory space contention. We cannot maximize per-core execution performance and HBM bandwidth utilization at the same time, due to on-chip memory space contention. As shown by in Figure 2, each core reserves an execution space for the currently executing operator and a preload space for the preloaded operators. To speed up per-core execution, a larger execution space is required (see §3.1 and Figure 5). To prevent HBM underutilization, a larger preload space is required (see §3.2 and Figure 6). With limited on-chip memory, we cannot expand both spaces.
Interconnect bandwidth contention. We cannot maximize HBM bandwidth utilization and minimize inter-core data sharing overhead at the same time, due to the interconnect bandwidth contention. As shown by in Figure 2, the on-chip interconnect carries both core-to-core traffic for inter-core data sharing and HBM controller-to-core traffic for preloading. When both traffic flows are heavy, the interconnect will be congested (see §3.3 and Figure 8).
Memory access contention. We cannot maximize per-core execution performance and minimize inter-core data sharing overhead at the same time, due to the memory access contention. As shown by in Figure 2, each core’s local memory is simultaneously accessed by the core itself for computing a tile, and by other cores for inter-core data sharing. For example on IPU, each core reads its local memory at full speed (128 bits/cycle (Graphcore, 2022)) when executing DL operators like MatMul, any other accesses will pause the execution. Upon contention, tile execution on this core reads data from local memory at slower speed, or even pauses entirely. The remote cores may also suffer from degraded SRAM bandwidth.
@@ 3. Performance Tradeoffs in Elk
As discussed in $\S$2.3, to efficiently use ICCA chips with HBM, we must trade-off multiple performance factors. We summarize how each performance factor is mapped to a compiler decision in Figure 4. First, increasing per-core execution space enables faster per-core execution with a larger tile size (§3.1). Second, increasing number of preloaded operators can better overlap on-chip execution and off-chip HBM load, improving HBM bandwidth utilization (§3.2). Third, increasing preload space for each preloaded operator reduces the inter-core data exchange overhead and the memory access contention, since the shared data can be duplicated on cores in advance to reduce the overhead of on-demand accesses to other cores (§3.3). We validate the insights with experiments on our ICCA chip emulator (see implementation in $\S$5) as follows.
@@ 3.1. Larger Execution Space Enables Faster Per-core Execution
Existing compilers focus on achieving higher on-chip execution performance using a given execution space size. However, they cannot find a proper execution space size by arbitrating the memory space contention in §2.3. Moreover, operators from the same model have diverse memory vs. time correlations. Thus, we also need to adjust the execution space size based on each operator’s performance characteristic, rather than allocating a fixed-sized execution space throughout the model execution.
@@ 3.2. Preloading More Operators Improves HBM Bandwidth Utilization
Figure 8. The total per-core interconnect bandwidth demand.
@@ 3.3. Larger Per-Operator Preload Space Reduces Inter-core Data Access Volume
Figure 7 shows that expanding preload space reduces the inter-core bandwidth demand. For each operator, we pick the fastest execution plan that fits in a given execution space size [^1] For each run, we use the optimal execution space size that gives the smallest total inference latency. See the description of the Static setup in §6.1.. MinPreload lets each core access all shared data from other cores at execution time, which requires the minimum preload space. MaxPreload lets HBM controllers broadcast as much shared data as possible at preload time, which requires the largest preload space. We profile the inter-core bandwidth demand ($\frac{\text{inter-core transfer volume}}{\text{per-core execution time}}$) of each core. MaxPreload significantly reduces the inter-core traffic.
[…]
@@ 4.2. Two-Level Inductive Operator Scheduling
The base case of induction is trivial, as the last operator has no succeeding operators to preload (i.e., preload number is always 0). For the inductive step, we show an example in Figure 10. In Figure 10 (a), Elk has finished scheduling all operators after Op5. Then, Elk schedules the execution of Op5 in Figure 10 (b). Elk enumerates all possible preload numbers for Op5. For each preload number, Elk invokes the cost-aware on-chip memory allocation algorithm ($\S$4.3) to determine the execution/preload space sizes for the involved operators and the estimated execution time of Op5.
For example, preload number 0 means we do not overlap Op5’s execution with any preload. The execution time of Op5 is minimized, but the overall execution time from Op5 to the end of the model is sub-optimal. For preload numbers 1 and 2, Elk overlaps the execution of Op5 with the preload of Op6, or the preloads of both Op6 and Op7. Though Op5’s execution time is longer, the overall execution times are better than preload number 0. As preload number 1 yields the lowest current-to-end time, Elk selects it for Op5.
@@
Our algorithm has $O(KN)$ complexity as we iterate through $N$ operators with up to $K$ preload numbers per operator. The algorithm provably finds the end-to-end plan with the shortest total time, assuming it can obtain the optimal execution time for each preload number. Lemma 4.1 and Theorem 4.2 formalize the algorithm.
@@
Let $1\leqii$<<NN. Suppose we have minimized $T_{\textit{end}}-T^{i+1}_{\textit{{s}-{exe}}}$. Then, there exists a preload number $p$ whose $T^{i}_{\textit{{s}-{exe}}}$ minimizes $T_{end}-T^{i}_{\textit{{s}-{exe}}}$, or maximizes $T^{i}_{\textit{{s}-{exe}}}$. Specifically, we have $T^{i}_{\textit{{e}-{exe}}}=\min(T^{i+1}_{\textit{{s}-{exe}}},T^{i+p+1}_{\textit{{s}-{pre}}})$, and $T^{i}_{\textit{{s}-{exe}}}=T^{i}_{\textit{{e}-{exe}}}-L^{i}_{exe}$ where $L^{i}_{exe}$ is the execution time of operator $i$ derived by the cost-aware memory allocation algorithm in $\S$4.3.
Proof.
First, to prove $T^{i}_{\textit{{e}-{exe}}}=\min(T^{i+1}_{\textit{{s}-{exe}}},T^{i+p+1}_{\textit{{s}-{pre}}})$ for any preload number $p$, we have (1) Opii must finish execution before Op(i+1)(i+1) starts execution, e.g., $T^{i}_{\textit{{e}-{exe}}}\leq T^{i+1}_{\textit{{s}-{exe}}}$; and (2) Opii’s execution can be overlapped with the preload of the next $p$ operators, which implies Opii’s execution must finish before the preload of Op(i+p+1)(i+p+1), e.g., $T^{i}_{\textit{{e}-{exe}}}\leq T^{i+p+1}_{\textit{{s}-{pre}}}$. Next, we prove the existence of $T^{i}_{\textit{{s}-{exe}}}$ that minimizes $T_{end}-T^{i}_{\textit{{s}-{exe}}}$. Suppose by contradiction that $T_{end}-T^{i+1}_{\textit{{s}-{exe}}}$ is minimized but there is no $T^{i}_{\textit{{s}-{exe}}}$ that minimizes $T_{end}-T^{i}_{\textit{{s}-{exe}}}$. Since our inductive step explored all $T^{i}_{\textit{{s}-{exe}}}$ values by enumerating all possible preload numbers, the only possible case is that we must explore more preload numbers to find the global $\max(T^{i}_{\textit{{s}-{exe}}})$, or the global $\max(T^{i}_{\textit{{e}-{exe}}})$ is greater than $T^{i+1}_{\textit{{s}-{exe}}}$, which means $T^{i+1}_{\textit{{s}-{exe}}}$ can be larger. This is a contradiction since $T_{end}-T^{i+1}_{\textit{{s}-{exe}}}$ is already minimized, e.g., $T^{i+1}_{\textit{{s}-{exe}}}$ is already maximized. ∎
@@ 4.3. Cost-Aware On-chip Memory Allocation
First, there are two types of intra-operator tradeoffs: (1) For the currently executing operator, Elk trades memory space for execution time ($\S$3.1). (2) For each preloaded operator, Elk trades memory space for this operator’s inter-core data exchange overhead ($\S$3.3). Second, there is an inter-operator tradeoff: as operators have different memory-time tradeoffs, we allocate more memory to operators that benefit more from a larger execution/preload space.
@@
Cost model for execution time. As DL workloads have predictable execution patterns (Feng et al., 2023; Miao et al., 2022; Zhu et al., 2022; Liu et al., 2024; Agrawal et al., 2024; Xue and Huang, 2025), Elk uses an accurate cost model to quickly estimate the performance of per-core execution and inter-core transfer. For each operator type (e.g., MatMul), we randomly generate tiles with varied shapes, and run each tile using one core on the target device. Then, we fit a linear tree model (Cerliani, 2022) using the tile shapes as inputs and the profiled execution times as outputs. For inter-core transfer, we fit a model for each network link using transfer volumes as inputs and transfer times as outputs. For each partition plan, Elk determines the tile-to-core mapping and orchestrates the inter-core transfer (e.g., the source/destination cores and intermediate hops of each transfer, see $\S$5). Elk uses the per-link cost model and the communication pattern to estimate the total transfer time. Figure 12 shows that Elk can accurately predict the execution and transfer times of an IPU chip. Elk can use different cost models (Miao et al., 2022; Zhu et al., 2022; Liu et al., 2024; Agrawal et al., 2024) for different hardware platforms.
Intra-operator tradeoff for preloading (Tradeoff 2 and 3 in Figure 11). For each preloaded operator, its partition plan is already decided in a previous step of the inductive operator scheduling ($\S$4.2). This execute-state plan is chosen for execution speed, which may use more memory space. As the operator is not currently executing, Elk assigns a memory-efficient preload-state plan. To start execution, a data distribution phase transforms the operator from preload- to execute-state by distributing the required data via the interconnect (e.g., Figure 3 (c)). It saves this operator’s preload space at the cost of extra inter-core data exchange overhead, compared to broadcasting the required data at preload time following the execute-state plan (e.g., Figure 3 (b)).
Each execute-state plan may have many preload-state plans, by configuring how much data is broadcasted on preload. On preload, if 4 cores share a data piece, we can evenly split it into 1, 2, or 4 chunks, and broadcast each chunk to 4, 2, or 1 cores. Each core receives $1$, $\frac{1}{2}$, or $\frac{1}{4}$ of the data on preload (this decides preload space size), and fetches the rest 0, $\frac{1}{2}$, or $\frac{3}{4}$ on data distribution. Elk finds the Pareto-optimal preload-state plans of each preloaded operator, by estimating their preload space sizes and data distribution times.
@@
For each search step, Elk examines the next plan with a smaller memory footprint along the Pareto-optimal curve for each operator. Elk selects the most “cost-effective” operator whose next plan has the largest ratio $\Delta=\frac{\text{reduced space size}}{\text{increased time}}$ compared to the currently selected plan. For example, in Figure 11, Op5 is the executing operator. Op6 and Op7 are preloaded operators. Elk updates the current plan for Op7 and proceed to the next search step. Elk stops when the total memory requirement does not exceed the available capacity.
In the worst case, Elk needs to examine all Pareto-optimal plans for all operators. Hence, it has $O(PK)$-time complexity for $K$ operators to fit on-chip and $P$ plans per operator. Combined with $\S$4.2, the complexity is $O(PK^{2}N)$ for $N$ operators ($K$ is also the number of possible preload numbers of each operator).